# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### v2

@@ 4.2 Training efficiency
In the following, we analyze the training time and peak GPU memory of training an early-exit model with pipeline parallelism, and propose some performance optimizations. We refer readers to the literature (e.g. the Megatron-LM series [67, 49, 36]) for thorough study of training efficiency with 3D parallelism for standard LLMs without early exits; we will use that as a baseline, and focus on the additional computational overhead caused by early exits.
@@
  • Implicit bubbles, i.e. dark gray areas in Figure 3(a). This is caused by load imbalance across pipeline stages, even though Transformer layers are evenly divided into stages in Megatron-LM [^3] For simplicity, we do not consider other flexible ways of dividing a LLM into pipeline stages, as dividing Transformer layers evenly remains one of the most practical and robust option in practice. We also do not consider the case where the input/output embedding matrix occupies a separate pipeline stage, although our techniques for training an early-exit LLM can be generalized to this case in principle.. In particular, the first stage has the additional computation for the input embedding layer, and more importantly, the last pipeline stage has the additional computation for the output logits (via the output embedding layer) as well as the training loss. For LLMs, these additional computational costs are not negligible, primarily due to large vocabulary sizes.
@@
  • Idle memory. Memory usage is also imbalanced across pipeline stages: earlier stages in the 1F1B schedule have to save the intermediate activations for more microbatches, and the first/last stage has to save an extra input/output embedding layer, plus the corresponding gradients and optimizer states. As a result, the first stage is typically the bottleneck of peak memory usage, while later stages have idle memory [36, Section 4.2.3].
Figure 3: One iteration of the 1F1B pipeline schedule in different cases. Within one iteration, each stage goes through a warm-up phase (forward steps of the beginning microbatches), a steady 1F1B (one-forward-one-backward) phase, and a cool-down phase (backward steps of the final microbatches). “Backbone forward/backward” stands for computation of Transformer layers on the backbone, while “Exit forward/backward” stands for computation of early-exit or final-exit layers. The number in each block denotes the index of the corresponding microbatch. Critical paths are marked by dashed red lines. In this figure, the degree of pipeline parallelism is set to 4, and each data batch is divided into 6 microbatches. It is assumed that for each stage, the ratio of time between forward computation for the backbone and forward computation for the early/final-exit layer is 2:1; moreover, we adopt the common assumption that the ratio of time between forward and backward computation is 1:2. Note that our analytical study does not rely on the specific values of these ratios. For clarity, we ignore computation related to the input embedding layer, as well as P2P communication latency between pipeline stages.
Remark 1.
@@
In our analysis, we assume no activation recomputation [10, 36] for simplicity. For clarity, we also ignore the P2P communication latency between pipeline stages, which is generally not the major concern for efficiency of pipeline parallelism.
4.2.1 Utilization of idle resources
@@
At first glance, one might expect that adding early exits to an LLM will incur a training overhead, in terms of time and memory, that is (at least) proportional to the number of additional model parameters. Fortunately, this is not the case for training with pipeline parallelism, based on the above analysis of idle resources in the 1F1B pipeline schedule. Indeed, adding one minimalistic early-exit layer (which, as defined in Section 2.2, has the same structure as the final output layer) to some middle (i.e. not the first or last) stage will only make its model size and theoretical forward/backward time match exactly those of the last stage. Therefore, the aforementioned implicit bubbles and some of the idle memory can be automatically utilized for computation related to the early-exit layers, leading to more balanced load across pipeline stages.
@@
More specifically, the overhead to training time caused by additional early exits can be negligible. If we choose $k$ middle stages and add one minimalistic early-exit layer to each of them, then the training time per iteration, i.e. time for processing one data batch, will (in theory) increase only by $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$, where $f_{\mathsf{EE}}$ and $b_{\mathsf{EE}}$ represent the time needed for one forward and backward pass of one microbatch for one minimalistic early-exit layer, respectively. To see this, first notice that the computation of early-exit layers in the steady 1F1B phase can be perfectly fit into the implicit bubbles. Therefore, the critical path remains the same as in the case without early exits, which consists of (1) the forward steps on all stages for the first microbatch, (2) the 1F1B steady phase on the last stage, and (3) the backward steps on all stages for the last microbatch. The early-exit layers, located separately on $k$ middle stages, will cause a $k\times f_{\mathsf{EE}}$ overhead to the first part of the critical path, and a $k\times b_{\mathsf{EE}}$ overhead to the third part, leading to the aforementioned claim on the overhead to training time. See Figures 3(a) and (b) for a visualization.
4.2.2 Further performance optimizations
@@
So far, we have analyzed training time and memory usage by model parameters. Another major component of memory usage, namely the memory for gradients and optimizer states, can be bounded by the memory for model parameters multiplied by a universal constant. One remaining issue that we have not addressed is the activation memory overhead due to early-exit computation. Most notably, the early-exit logits for one microbatch have size $s\times b\times V$, where $s$ is the maximum sequence length, $b$ is the microbatch size, and $V$ is the vocabulary size. If the $i$-th stage has one early exit, then a vanilla implementation of training with early exits using the 1F1B pipeline schedule (as shown in Figure 3(b)) will cause a significant memory overhead of size $s\times b\times V\times(P-i+1)$, where $P-i+1$ is the number of in-flight microbatches [36] for Stage $i$.
@@
Remark 2.
One can check that, with the above adjustment, our analysis in Section 4.2.1 about the training time overhead caused by early exits remains valid after minor modifications. More specifically, the time overhead of one training iteration is still $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$; the only difference is that this whole overhead comes from the backward steps of the last microbatch, i.e. the third part of the critical path in Figure 3(c). One can further reduce this overhead to $k\times b_{\mathsf{EE}}$ by moving the forward pass of the early-exit layer on each stage for each cool-down microbatch to the explicit bubble in front of the corresponding backward step (i.e. before communication with the next stage). With that said, our implementation does not include this modification, as it brings limited gains at the cost of complicating our codebase.
@@
  • Avoid adding too many early exits to the LLM. Despite higher flexibility during inference, the gain of adding many early exits (e.g. one per layer) might be marginal, and comes at the cost of excessive overhead for training and inference, which is especially the case for LLMs due to large vocabulary sizes. Similar observations and advice have been made recently by the authors of [3] as well.
@@
Remark 3.
Recall from Section 3.1 that, with EE-LLM, users can choose more expressive and powerful early-exit layers beyond the minimalistic structure. Similarly, more than one early exit can be added to each pipeline stage, which provides more flexible choices of exits during inference. These benefits, of course, come at the cost of higher overhead for training, and potentially for inference as well444 There is no clear answer to whether additional modules at early-exit layers will improve or hurt the overall inference speed. There is certainly higher overhead for the computation of each early-exit layer; on the other hand, higher flexibility and adaptivity of the early exits can potentially enable them to produce better outputs and get selected more often during inference, leading to overall faster generation of a complete sequence. For similar reasons, there is no clear positive or negative correlation between the number of early exits and the overall speed of generating a sequence.; with EE-LLM, users can conveniently choose the most suitable configurations for their own use cases. We refer interested readers to Appendix A for formal analysis of training efficiency in these general cases.
Numerical examples.
Figure 4: The forward time, backward time, and peak GPU memory of each pipeline stage for a standard 7B GPT Transformer, as well as its early-exit version that has one minimalistic early-exit layer (without layer normalization) added to each middle stage. Degrees of pipeline, tensor and data parallelism are 4, 1, and 1, respectively; the microbatch size is 2, global batch size is 128, and sequence length is 2048. Note that the forward computation of early-exit layers has been deferred to the backward steps, hence not included in “forward time” of the first plot, but in “backward time” of the second plot.

### v3

@@ Appendix A Training efficiency
This section is an extended version of Section 3.2. In the following, we analyze the training time and peak GPU memory of training an early-exit model with pipeline parallelism, and propose some performance optimizations. We refer readers to the literature (e.g. the Megatron-LM series (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022)) for thorough study of training efficiency with 3D parallelism for standard LLMs without early exits; we will use that as a baseline, and focus on the additional computational overhead caused by early exits.
@@
  • Implicit bubbles, i.e. dark gray areas in Figure 3(a). This is caused by load imbalance across pipeline stages, even though Transformer layers are evenly divided into stages in Megatron-LM [^3] For simplicity, we do not consider other flexible ways of dividing an LLM into pipeline stages, as dividing Transformer layers evenly remains one of the most practical and robust options in practice. We also do not consider the case where the input/output embedding matrix occupies a separate pipeline stage, although our techniques for training an early-exit LLM can be generalized to this case in principle.. In particular, the first stage has the additional computation for the input embedding layer, and more importantly, the last pipeline stage has the additional computation for the output logits (via the output embedding layer) as well as the training loss. For LLMs, these additional computational costs are not negligible, primarily due to large vocabulary sizes.
@@
  • Idle memory. Memory usage is also unbalanced across pipeline stages: earlier stages in the 1F1B schedule have to save the intermediate activations for more microbatches, and the first/last stage has to save an extra input/output embedding layer, plus the corresponding gradients and optimizer states. As a result, the first stage is typically the bottleneck of peak memory usage, while later stages have idle memory (Korthikanti et al., 2022).
Remark A.1.
@@
In our analysis, we assume no activation recomputation (Chen et al., 2016; Korthikanti et al., 2022) for simplicity. For clarity, we also ignore the P2P communication latency between pipeline stages, which is generally not the major concern for the efficiency of pipeline parallelism.
A.1 Utilization of idle resources
@@
At first glance, one might expect that adding early exits to an LLM will incur a training overhead, in terms of time and memory, that is (at least) proportional to the number of additional model parameters. Fortunately, this is not the case for training with pipeline parallelism, based on the above analysis of idle resources in the 1F1B pipeline schedule. Indeed, adding one minimalistic early-exit layer (which has the same structure as the final output layer) to some middle (i.e. not the first or last) stage will only make its model size and theoretical forward/backward time match exactly those of the last stage. Therefore, the aforementioned implicit bubbles and some of the idle memory can be automatically utilized for computation related to the early-exit layers, leading to more balanced load across pipeline stages.
@@
More specifically, the overhead to training time caused by additional early exits can be negligible. If we choose $k$ middle stages and add one minimalistic early-exit layer to each of them, then the training time per iteration, i.e. time for processing one data batch, will (in theory) increase only by $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$, where $f_{\mathsf{EE}}$ and $b_{\mathsf{EE}}$ represent the time needed for one forward and backward pass of one microbatch for one minimalistic early-exit layer, respectively. To see this, first notice that the computation of early-exit layers in the steady 1F1B phase can perfectly fit into the implicit bubbles. Therefore, the critical path remains the same as in the case without early exits, which consists of (1) the forward steps on all stages for the first microbatch, (2) the 1F1B steady phase on the last stage, and (3) the backward steps on all stages for the last microbatch. The early-exit layers, located separately on $k$ middle stages, will cause a $k\times f_{\mathsf{EE}}$ overhead to the first part of the critical path, and a $k\times b_{\mathsf{EE}}$ overhead to the third part, leading to the aforementioned claim on the overhead to training time. See Figures 3(a) and (b) for a visualization.
A.2 Further performance optimizations
@@
So far, we have analyzed training time and memory usage by model parameters. Another major component of memory usage, namely the memory for gradients and optimizer states, can be bounded by the memory for model parameters multiplied by a universal constant. One remaining issue that we have not addressed is the activation memory overhead due to early-exit computation. Most notably, the early-exit logits for one microbatch have size $s\times b\times V$, where $s$ is the maximum sequence length, $b$ is the microbatch size, and $V$ is the vocabulary size. If the $i$-th stage has one early exit, then a vanilla implementation of training with early exits using the 1F1B pipeline schedule (as shown in Figure 3(b)) will cause a significant memory overhead of size $s\times b\times V\times(P-i+1)$, where $P-i+1$ is the number of in-flight microbatches (Korthikanti et al., 2022) for Stage $i$.
@@
Remark A.2.
One can check that, with the above adjustment, our analysis in Section A.1 about the training time overhead caused by early exits remains valid after minor modifications. More specifically, the time overhead of one training iteration is still $k\times(f_{\mathsf{EE}}+b_{\mathsf{EE}})$; the only difference is that this whole overhead comes from the backward steps of the last microbatch, i.e. the third part of the critical path in Figure 3(c). One can further reduce this overhead to $k\times b_{\mathsf{EE}}$ by moving the forward pass of the early-exit layer on each stage for each cool-down microbatch to the explicit bubble in front of the corresponding backward step (i.e. before communication with the next stage). With that said, our implementation does not include this modification, as it brings limited gains at the cost of complicating our codebase.
@@
  • Avoid adding too many early exits to the LLM. Despite higher flexibility during inference, the gain of adding many early exits (e.g. one per layer) might be marginal, and comes at the cost of excessive overhead for training and inference, which is especially the case for LLMs due to large vocabulary sizes. Similar observations and advice have been made recently by the authors of (Bae et al., 2023) as well.
@@
Remark A.3.
Recall from Section 2 that, with EE-LLM, users can choose more expressive and powerful early-exit layers beyond the minimalistic structure. Similarly, more than one early exit can be added to each pipeline stage, which provides more flexible choices of exits during inference. These benefits, of course, come at the cost of a higher overhead for training, and potentially for inference as well444 There is no clear answer to whether additional modules at early-exit layers will improve or hurt the overall inference speed. There is certainly a higher overhead for the computation of each early-exit layer; on the other hand, higher flexibility and adaptivity of the early exits can potentially enable them to produce better outputs and get selected more often during inference, leading to overall faster generation of a complete sequence. For similar reasons, there is no clear positive or negative correlation between the number of early exits and the overall speed of generating a sequence.; with EE-LLM, users can conveniently choose the most suitable configurations for their own use cases. A formal analysis of training efficiency in these general cases will soon be presented in Appendix A.3.
Figure 9: The forward time, backward time, and peak GPU memory of each pipeline stage for a standard 7B GPT Transformer, as well as its early-exit version that has one minimalistic early-exit layer (without layer normalization) added to each middle stage. Degrees of pipeline, tensor and data parallelism are 4, 1, and 1, respectively; the microbatch size is 2, the global batch size is 128, and the sequence length is 2048. Note that the forward computation of early-exit layers has been deferred to the backward steps, hence not included in “forward time” of the first plot, but in “backward time” of the second plot.